# Win Ratio with Multiple Thresholds for Composite Endpoints

### v4

@@ Abstract
Composite endpoints consisting of both terminal and non-terminal events, such as death and hospitalization, are frequently used as primary endpoints in cardiovascular clinical trials. The Win Ratio method (WR) employs a hierarchical structure to combine fatal and non-fatal events by giving death information an absolute priority, which can adversely affect power if the treatment effect is mainly on the non-fatal outcomes. We hereby propose the Win Ratio with Multiple Thresholds (WR-MT) that releases the strict hierarchical structure of the standard WR by adding stages with non-zero thresholds. A weighted adaptive approach is also developed to determine the thresholds in WR-MT. This method preserves the statistical properties of the standard WR but can sometimes increase the chance to detect treatment effects on non-fatal events. We show that WR-MT has a particularly favorable performance than standard WR when the second layer has stronger signals and otherwise comparable performance in our simulations that vary the follow-up time, the correlation between events, and the treatment effect sizes. A case study based on the Digitalis Investigation Group clinical trial data is presented to further illustrate our proposed method. An R package “WRMT” that implements the proposed methodology has been developed.
Keywords Clinical trials  $\cdot$ Composite Survival Endpoints  $\cdot$ Endpoint Hierarchy  $\cdot$ Pairwise comparison  $\cdot$ Win ratio  $\cdot$ Win statistics
@@ 1 Introduction
Composite time-to-event endpoints are frequently employed in cardiovascular and other clinical trials that evaluate the impact of new candidate treatments, to possibly improve statistical efficiency over a single endpoint and to dispense the need for multiplicity adjustment with multiple endpoints (Mao and Kim, 2021). For example, in a standard scenario, two common events used to define the composite endpoint are survival time and time to hospitalization. Such choices arise naturally since cardiovascular diseases have high morbidity and are more prevalent in elderly populations, where death is an immediate concern and a likely cause for informative censoring. The traditional analysis focuses on the time to the first event (either death or hospitalization). However, the time-to-first event approach often fails to consider the differential clinical importance between death and hospitalization, and does not have the capacity to address subsequent occurrence of a second component endpoint (e.g., a death event after a hospitalization event).
To address the potential limitations of the time-to-first event approach, Pocock et al. (2012) proposed the matched and unmatched win ratio method. For the unmatched setting, hypothesis testing is conducted using the method introduced by Finkelstein and Schoenfeld (1999). We use the abbreviation WR to refer to either the win ratio method or the win ratio measure. The full term “win ratio” will be used for the win ratio measure when there is ambiguity. WR and its related methods have quickly gained popularity in the past several years and a few software has been developed to date. An increasing number of clinical trials have also considered WR as one of the mainstream approaches for addressing composite outcomes (Gasparyan et al., 2021). As an example, the EMPULSE trial (ClinicalTrials.gov identifier: NCT0415775), the VIP-ACS trial (NCT04001504), and the DAPA-HF trial (NCT03036124) included WR as the primary outcome measure in the protocol. As discussed by Finkelstein and Schoenfeld (1999), another potential advantage of prioritized endpoints is to offer a systematic approach to combine the event time outcome with other types of clinical outcomes, such as longitudinal biomarkers, particularly when event outcomes necessitate a longer time to accumulate. Examples include the development of AIDS and longitudinal changes in viral load in HIV studies, and the onset of end-stage renal disease and changes in estimated glomerular filtration rate in kidney studies (Fergusson et al., 2018).
As the application of WR continues to expand in practical settings, there has been a growing focus towards its methodological research. The stratified WR was introduced to adjust for discrete baseline variables (Dong et al., 2018; Gasparyan et al., 2021). Methods for statistical inference (e.g, in constructing confidence intervals and conducting hypothesis testing) with the WR measure were also proposed and extensively investigated (Luo et al., 2015; Bebu and Lachin, 2016; Dong et al., 2016; Mao, 2019). Oakes (2016) discussed the estimation of WR for censored observations. In the presence of potentially covariate-dependent right censoring, the inverse-probability-of-censoring weighted WR was proposed to eliminate the dependence of the effect measure on the censoring distribution (Dong et al., 2020, 2021). In the broader family of methods to which WR belongs, known as win statistics (Dong et al., 2023; Verbeeck et al., 2023), several variations with similar outcome prioritization are available, including but not limited to the win loss (Luo et al., 2017), the generalized pairwise comparisons (corresponding to the effect measure of net benefit) (Buyse, 2010), the win probability (Gasparyan et al., 2021), the win odds that address ties (Brunner et al., 2021), and the event-specific WR (Yang and Troendle, 2021; Yang et al., 2022). See also Péron et al. (2018) and Deltuvaite-Thomas et al. (2023) for discussions on the censoring issue in generalized pairwise comparisons and additional extensions. More recently, regression methods that directly explore the association between baseline covariates and win functions have also been developed (Mao and Wang, 2021; Wang and Mao, 2022; Song et al., 2023).
Despite its popularity, the standard WR is subject to possible limitations. As pointed out by Redfors et al. (2020), the use of WR may cause a substantial loss of power when the treatment does not affect the top layer of the hierarchy and the censoring rate of this top-layer outcome is low. As a concrete example, we let survival time and time to hospitalization form the first and second layers of WR, and assume the treatment only affects the time-to-hospitalization layer. If the death rate is high, the majority of pairwise comparison outcomes are determined by the survival times. This leaves limited opportunity for moving onto the time-to-hospitalization layer to detect the true treatment signal. More broadly, this phenomenon, mostly due to the strict hierarchical structure of WR, can occur if any higher layer fails to show a treatment signal despite strong treatment signals that may exist the lower layers. In other words, comparison results on the higher layer may conceal the true treatment effect on the lower layer, because no or little comparison will be made on the lower layer if few ties are present in the higher layer. Recognizing this potential limitation of the standard WR, we introduce multiple thresholds to the definition of the win function, and refer to this new approach as the Win Ratio with Multiple Thresholds (WR-MT). Specifically, our win function is based on pairwise comparisons that employ multiple pre-specified thresholds applied sequentially, such that a definitive win (or loss) is declared only when the observed difference exceeds these thresholds. Ultimately, the thresholds are shrunk to zero, enabling the win determination to reflect the continuous nature of the outcome (Redfors et al., 2020). This is in contrast to the definition of the standard win functions, for which a single threshold (often zero) is typically defined for each endpoint.
@@
The general concept of looking at multiple thresholds in pairwise comparisons dates back to Buyse (2010). Specifically, under the framework of generalized pairwise comparisons, they considered multiple comparison thresholds for a continuous or time-to-event endpoint, for which the successive thresholds were designed to reflect different levels of clinical differences. In follow-up work (Péron et al., 2016; Péron et al., 2019), multiple thresholds were also considered separately in individual tests for sensitivity analysis. In general, the implications of using multiple thresholds within a single test have not been fully elucidated, nor there exist general operational guidelines when multiple thresholds are considered in a single test. Moreover, no prior work has integrated multiple thresholds and multiple outcomes within a unified testing framework; hence our proposed WR-MT method fills this critical gap. In our setting with multiple outcomes, the motivation for using multiple thresholds in WR-MT is to provide lower layers higher opportunities to contribute information to the win function when necessary, without compromising the priority of higher layers. Finally, in addition to formalizing the architecture with multiple thresholds for multiple endpoints, we also introduce a data-driven weighted adaptive approach to automatically determine possible thresholds, offering some convenience in specifying the thresholds without affecting the test size.
@@
To summarize, as a modified version of WR, WR-MT maintains the statistical properties of WR but relaxes the strict hierarchical structure of WR with intuitive modifications to the comparison process. At each layer, WR-MT retains all the ties in WR but could lead to different results without ties. By potentially correcting the results of some pairwise comparisons, WR-MT may increase statistical power, especially when the treatment effect signal is weak on the higher layer. The design of the weighted adaptive thresholds is capable of detecting the treatment effects on the higher layer while reducing the possible noise (a detailed mechanism will be illustrated later). To better demonstrate the advantage of WR-MT, we compare the performance of the proposed WR-MT approach versus WR via simulations. In our simulations, three critical influencing factors, namely the maximal follow-up time, the correlation between two endpoints, and the treatment effect size, are considered. For further illustration, we apply both methods to reanalyze data from the Digitalis Investigation Group (DIG) clinical trial (NCT00000476). Throughout the paper, we focus on the basic setting of two-component endpoints and two thresholds per endpoint to facilitate the exposition of the key idea. We also provide a set of generalized formulas with an arbitrary finite number of component endpoints and additional thresholds for completeness.
To reiterate, our main contributions are three-folded. First, we formalize the multiple-threshold structure for WR with multiple endpoints, where thresholds are applied alternatively across endpoints. This approach relaxes strict prioritization while preserving the hierarchical ordering inherent in the WR, and can be readily employed in other win statistics and related methodology extensions. Second, we design a weighted adaptive procedure with interpretable tuning parameters to determine the thresholds. This approach offers convenience in specifying the thresholds in the absence of strong clinical priors. Third, we conduct an extensive set of simulations and report on a case study to systematically evaluate our proposed method and address how correlations among endpoints may influence the power of WR and WR-MT. The remainder of the paper is organized as follows. In Section 2, we describe WR-MT, explain its key features and potential advantages, and generalize WR-MT to multiple endpoints. Section 3 reports on a simulation study that examines the operating characteristics of WR-MT. In Section 4, the proposed WR-MT is applied to analyze the DIG trial. Section 5 concludes with a discussion.
@@ 2 Statistical Methods
In this section, we first review the standard WR method, then we propose our new WR-MT under a simple yet common setting with two component time-to-event endpoints and two thresholds per endpoint, followed by the weighted adaptive approach used to determine thresholds. A general formulation of WR-MT with multiple component endpoints and an arbitrary number of thresholds will be provided in Appendix A. Starting with the simple setting, we assume there are two layers, where the first layer is survival time and the second layer is time to hospitalization, which is a typical non-fatal event of clinical importance. To fix the notation, we assume a clinical trial with $n$ participants and two treatment arms with $Z_{i}=0$ if the participant is in the control group and $Z_{i}=1$ if in the treatment group. For the $i$-th participant, let $D_{i}$ be the (observed) survival time; $\Delta_{Di}$ denote the censoring indicator for $D_{i}$ such that $\Delta_{Di}=0$ if death event is observed and $\Delta_{Di}=1$ if the participant is right censored. We write $T_{1i}$ to denote the observed time to hospitalization, and similarly $\Delta_{1i}$ to denote the censoring indicator for this non-fatal event. Throughout, we let $I(\cdot)$ be the indicator function and $\text{Sign}(\cdot)$ be the sign function.
2.1 Standard win ratio approach: A review
@@
WR is based on pairwise comparisons across all study participants. For the comparison between participants $i$ and $j$, we define the win-loss score $U_{ij}$ to be 1 if $i$ has a more favorable outcome (win) than $j$, 1-1 if $i$ has a less favorable outcome (loss) than $j$, and 0 if neither is more favorable hence the comparison is uninformative or indeterminate (tie). Within each pairwise comparison, for the calculation of $U_{ij}$, one first examines the survival time to assess if one participant lives longer than the other. If $i$ and $j$ have the same days alive, or the comparison is uninformative due to censoring, the time to hospitalization is then compared to assess if one participant has experienced a longer time before hospitalization than the other. If the time to hospitalization is the same, a tie will be concluded for the comparison between participants $i$ and $j$. WR of the treatment group is then calculated by
$R=\frac{\sum_{i=1}^{n}Z_{i}\sum_{j=1}^{n}(1-Z_{j})I(U_{ij}=1)}{\sum_{i=1}^{n}Z_{i}\sum_{j=1}^{n}(1-Z_{j})I(U_{ij}=-1)}.$
Thus, $R>1$ means the treatment group has more wins than losses in pairwise comparisons (i.e., treatment is potentially better than control). Hypothesis testing can then be formulated to make statistical inference about the treatment effect, drawing results from the theory of U-statistics (Bebu and Lachin, 2016) (see details of computing variance and constructing hypothesis test in Finkelstein and Schoenfeld (1999), Pocock et al. (2012), and Luo et al. (2015)). As a brief recap, if the only interest is to test the significance of the treatment effect without constructing the confidence interval of WR itself, the test statistic $S=\sum_{i=1}^{n}Z_{i}\sum_{j=1}^{n}U_{ij}$ can be employed. It has been shown that the test statistic (Finkelstein and Schoenfeld, 1999), $S$, follows an asymptotic normal distribution with mean zero under the null and has a closed-form variance estimator given by
@@
2.2 Win ratio with multiple thresholds
In order to release the strict hierarchical structure of WR and allow for a controllable degree of priority that survival time has against time to hospitalization, we propose the win ratio with multiple thresholds, termed WR-MT. With WR-MT, one first compares two participants with larger thresholds through the hierarchical structure before moving to the set of comparisons based on lower thresholds. The idea of WR-MT is illustrated in Figure 1, where the first column of comparisons describes the comparisons under the larger thresholds $d$ and $t_{1}$, and the second column of comparisons describes the comparisons using 0 as the threshold. One will note that the standard WR is a special case of WR-MT that only involves the second column of comparisons for calculating the win-loss score.

### v5

# Generalizing the Finkelstein–Schoenfeld Test to Incorporate Multiple Alternating Thresholds

@@ Abstract
Composite endpoints consisting of both terminal and non-terminal events, such as death and hospitalization, are frequently used in cardiovascular clinical trials. The Finkelstein-Schoenfeld (FS) test provides a way to employ a hierarchical structure to combine fatal and non-fatal events by giving death information an absolute priority, which may limit the contribution of clinically meaningful non-fatal events. To provide a more flexible alternative, we propose the Finkelstein–Schoenfeld with Multiple Thresholds (FS-MT) test, which extends the standard FS test by incorporating multiple thresholds applied sequentially and alternating across endpoints. A weighted adaptive approach is also developed to help determine the thresholds in FS-MT. The proposed approach retains the statistical properties of the FS test while allowing more flexible use of information from lower-priority events. We evaluate the operating characteristics of the proposed test through simulations that vary the follow-up time, the correlation between events, and the treatment effect sizes. A case study based on the Digitalis Investigation Group clinical trial data is presented to further illustrate our proposed method. An R package “FSMT” that implements the proposed methodology has been developed.
Keywords Clinical trials $\cdot$ Composite Survival Endpoints $\cdot$ Endpoint Hierarchy $\cdot$ Generalized Pairwise Comparisons $\cdot$ Win ratio $\cdot$ Win statistics
@@ 1.1 Background and literature review
Composite time-to-event endpoints are frequently employed in cardiovascular and other clinical trials that evaluate the impact of new candidate treatments, to possibly improve statistical efficiency over a single endpoint and to dispense the need for multiplicity adjustment with multiple endpoints (mao2021statistical). For example, in a standard scenario, two common events used to define the composite endpoint are survival time and time to hospitalization. Such choices arise naturally since cardiovascular diseases have high morbidity and are more prevalent in elderly populations, where death is an immediate concern and a likely cause for informative censoring. The traditional analysis focuses on the time to the first event (either death or hospitalization). However, this approach often fails to consider the differential clinical importance between death and hospitalization, and does not have the capacity to address subsequent occurrence of a second component endpoint (e.g., a death event after a hospitalization event).
The FS test originally proposed by finkelstein1999combining offers an approach to address these potential limitations. This test relies on pairwise comparisons between participants’ outcomes using a hierarchical structure, in which the survival time endpoint is often prioritized over non-fatal events. This procedure aligns the analysis with clinical priorities, ensuring that more serious events receive higher attention in evaluating treatment effects. In addition, as discussed by finkelstein1999combining, another advantage of prioritized endpoints is to offer a systematic approach to combine the event time outcome with other types of clinical outcomes, such as longitudinal biomarkers, particularly when event outcomes necessitate a longer time to accumulate. This concept of hierarchical pairwise comparisons was later formalized into the net benefit (NB) treatment effect measure by buyse2010generalized, and was subsequently popularized in the cardiovascular research community through the introduction of the win ratio (WR) measure by pocock2012win. In recent years, the WR and the broader framework of Generalized Pairwise Comparisons (GPC) have gained substantial attention. Owing to their advantages, an increasing number of clinical trials have adopted hierarchical pairwise comparison methods for analyzing composite endpoints (gasparyan2021adjusted). For instance, the EMPULSE trial (ClinicalTrials.gov identifier: NCT0415775), the VIP-ACS trial (NCT04001504), and the DAPA-HF trial (NCT03036124) included WR as the primary outcome measure in their protocols. Notably, the ATTR-ACT trial (NCT01994889), which evaluated Tafamidis for transthyretin amyloid cardiomyopathy, employed the FS test as its primary hypothesis testing method, resulting in the first U.S. Food and Drug Administration (FDA) approval of a therapy based on a GPC framework in 2019.
As the application of the GPC framework continues to expand in practical settings, there has been a growing focus on methodological research for hierarchical pairwise comparison methods. As an example, methods for statistical inference (e.g, in constructing confidence intervals and conducting hypothesis testing) with the WR measure were extensively studied (luo2015alternative; bebu2016large; dong2016generalized; mao2019alternative). The stratified WR was introduced to adjust for discrete baseline variables (dong2018stratified; gasparyan2021adjusted). Beyond the FS test, NB and WR measures, the GPC framework now encompasses a rich family of related effect measures and methodological variations, including but not limited to the win–loss statistic (luo2017weighted), win probability (gasparyan2021adjusted), win odds (brunner2021win), and event-specific win ratio (yang2021eventct). Each of these measures shares a similar pairwise comparison foundation but offers different ways to summarize or weight the comparative information. There have also been methodological developments for the GPC framework in general, such as statistical inference methods (see ozenne2025inference for a review and comparison), stratification (dong2023stratified; backer2025stratification), handling of censoring (peron2018extension; deltuvaite2023generalized; deltuvaite-thomas2025rightcensored), and beyond. See dong2023win, verbeeck2023generalized, and buyse2025handbook for further details and discussions on the GPC framework.
1.2 Objectives and contributions
Despite its successful usage and growing popularity, testing the null hypothesis with the standard FS test (or other approximately equivalent hypothesis testing procedures within the GPC framework) may sometimes limit sensitivity to treatment effects on lower-priority endpoints. As discussed by redfors2020win, when the top layer dominates comparisons, such as when mortality is frequent and largely unaffected by treatment, potential signals at subsequent layers may be underrepresented, which is an inherent feature of comparison schemes with a strict hierarchical structure. To provide an alternative perspective under such settings, we propose the Finkelstein–Schoenfeld with Multiple Thresholds (FS-MT) test, which builds upon the FS framework while allowing more flexible use of information from lower-priority endpoints. Specifically, our win function is based on pairwise comparisons that employ multiple pre-specified thresholds applied sequentially and alternating across endpoints, such that a definitive win (or loss) is declared only when the observed difference exceeds these thresholds. Ultimately, the thresholds are shrunk to minimal clinically meaningful differences. This is in contrast to the definition of the standard win functions, for which a single threshold is typically defined for each endpoint. While such minimal clinically meaningful differences may be nonzero, for example, a difference of 5 points in the Kansas City Cardiomyopathy Questionnaire Overall Score is often considered clinically significant (voors2022sglt2), we use zero in our illustration for simplicity. Nonzero thresholds are naturally supported by both the standard FS test and the proposed FS-MT test.
@@ 1.2 Objectives and contributions
The general concept of looking at multiple thresholds in pairwise comparisons dates back to buyse2010generalized. Specifically, for the NB measure, they considered multiple comparison thresholds, for which the successive thresholds were designed to reflect different levels of clinical differences. However, their initial work did not explicitly explore the use of multiple thresholds in settings involving more than one endpoint, particularly not in scenarios where thresholds are to be applied in an alternating manner across different endpoints. In follow-up work (peron2016assessment; peron2019benefit), multiple thresholds were also considered separately in individual tests for sensitivity analysis. In general, the implications of using multiple thresholds for each endpoint within a single test have not been fully elucidated, nor there exist general operational guidelines when multiple thresholds for each endpoint are considered in a single test. Moreover, no prior work has integrated multiple thresholds and multiple outcomes within a unified testing framework; hence our proposed FS-MT method fills this gap. In our setting with multiple outcomes, the motivation for using multiple thresholds in FS-MT is to provide lower layers higher opportunities to contribute information to the win function when necessary. Finally, in addition to formalizing the architecture with multiple thresholds for multiple endpoints, we also introduce a data-driven weighted adaptive approach to automatically determine possible thresholds, offering some convenience in specifying the thresholds without affecting the test size.
@@
To summarize, as a modified version of FS test, the FS-MT test maintains the statistical properties of the standard FS test but relaxes its strict hierarchical structure with intuitive modifications to the comparison process. At each layer, the FS-MT test retains all the ties in the FS test but could lead to different results without ties. To better demonstrate the performance of the FS-MT test, we compare the proposed FS-MT test with the FS test via simulations. In our simulations, three critical influencing factors, namely the maximal follow-up time, the correlation between two endpoints, and the treatment effect size, are considered. For further illustration, we apply both methods to reanalyze data from the Digitalis Investigation Group (DIG) clinical trial (NCT00000476).
@@ 2 Statistical Methods
In this section, we first review the standard FS test and commonly used treatment effect measures in the GPC framework. Then we describe our proposed FS-MT test under a simple yet common setting with two component time-to-event endpoints and two thresholds per endpoint, followed by a weighted adaptive approach to determine thresholds. A general formulation of FS-MT with multiple component endpoints and an arbitrary number of thresholds will be provided in Supporting Information S1. Starting with the simple setting, we assume there are two layers, where the first layer is survival time and the second layer is time to hospitalization, which is a typical non-fatal event of clinical importance. We assume a clinical trial with $n$ participants and two treatment arms with $Z_{i}=0$ if the participant is in the control group and $Z_{i}=1$ if in the treatment group. For the $i$-th participant, let $D_{i}$ be the (observed) survival time, $\Delta_{Di}$ denote the censoring indicator for $D_{i}$ such that $\Delta_{Di}=0$ if death event is observed and $\Delta_{Di}=1$ if the death time is right censored. We write $T_{1i}$ to denote the observed time to hospitalization, and similarly $\Delta_{1i}$ to denote the censoring indicator; that is $\Delta_{Ti}=0$ if non-fatal event is observed and $\Delta_{Ti}=1$ if right censored. Throughout, we let $I(\cdot)$ be the indicator function and $\text{Sign}(\cdot)$ be the sign function.
2.1 Standard FS test: A review
@@
The FS test is based on pairwise comparisons across all study participants. For the comparison between participants $i$ and $j$, we define the win-loss score $U_{ij}$ to be 1 if $i$ has a more favorable outcome (win) than $j$, 1-1 if $i$ has a less favorable outcome (loss) than $j$, and 0 if neither is more favorable hence the comparison is uninformative or indeterminate (tie). Within each pairwise comparison, for the calculation of $U_{ij}$, one first examines the survival time to assess if one participant lives longer than the other. If participants $i$ and $j$ have the same days alive, or the comparison is uninformative due to censoring (hence tied on survival time), the time to hospitalization is then compared to assess if one participant has experienced a longer time before hospitalization than the other. If the time to hospitalization is the same or uninformative, a tie will be concluded for the comparison between participants $i$ and $j$.
To test the null hypothesis of no treatment effect with such composite endpoints, the FS test considers the following test statistic:
$S=\sum_{i=1}^{n}Z_{i}\sum_{j=1}^{n}U_{ij}.$
Here, $S$ is the sum of total win-loss scores (net win scores), obtained by comparing each treated participant against all participants. finkelstein1999combining have shown that, under the null, $S$ follows an asymptotic normal distribution with mean zero and admits a closed-form variance estimator given by
@@
The FS test can be viewed as a hypothesis testing method within the GPC framework (buyse2010generalized; buyse2025handbook). While the FS test focuses on hypothesis testing, the pairwise comparison results in the GPC framework can also be used to construct treatment effect measures that quantify both the magnitude and direction of treatment benefit for composite endpoints. Let $N_{+}=\sum_{i=1}^{n}Z_{i}\sum_{j=1}^{n}(1-Z_{j})I(U_{ij}=1)$, $N_{-}=\sum_{i=1}^{n}Z_{i}\sum_{j=1}^{n}(1-Z_{j})I(U_{ij}=-1)$, $N_{0}=\sum_{i=1}^{n}Z_{i}\sum_{j=1}^{n}(1-Z_{j})I(U_{ij}=0)$ denote the total numbers of wins, losses, and ties obtained by treatment participants when compared against control participants, respectively. It is immediate that the FS test statistic is given by $S=(N_{+}-N_{-})+\sum_{i=1}^{n}\sum_{j=1}^{n}Z_{i}Z_{j}U_{ij}$. Based on these quantities, treatment effect measures commonly used in the GPC framework can be calculated. For example, net benefit (NB), win odds (WO), and win ratio (WR) are obtained as:
$\text{NB}=\frac{N_{+}-N_{-}}{N_{+}+N_{-}+N_{0}},\qquad\text{WO}=\frac{N_{+}+0.5N_{0}}{N_{-}+0.5N_{0}},\qquad\text{WR}=\frac{N_{+}}{N_{-}}.$ (1)
For these three GPC measures, $\text{NB}=0$, $\text{WO}=1$, and $\text{WR}=1$ indicate no treatment effect, whereas $\text{NB}<0$, $\text{WO}<1$, and $\text{WR}<1$ suggest a potential negative treatment effect, and the opposite directions ($\text{NB}>0$, $\text{WO}>1$, $\text{WR}>1$) imply a positive treatment effect. These estimators are also referred to as win statistics, and methods for statistical inference have been developed based on each win statistic, leveraging the theory of U-statistics, randomization-based procedure, and beyond; see, for example, buyse2010generalized, luo2015alternative, bebu2016large, brunner2021win, and buyse2025handbook. In what follows, we focus primarily on hypothesis testing by introducing a new variant of the FS test that accommodates alternating thresholds across endpoints.
2.2 Alternating multiple thresholds across endpoints
In order to release the strict hierarchical structure of the FS test and allow for a controllable degree of priority that survival time has against time to hospitalization, we propose to alternate multiple thresholds across endpoints in the original FS test, termed FS-MT test. With the FS-MT test, one first compares two participants with larger thresholds through the hierarchical structure before moving to the set of comparisons based on lower thresholds. The idea of the FS-MT test is illustrated in Figure 1, where the first column of comparisons describes the comparisons under the larger thresholds $d$ and $t_{1}$, and the second column of comparisons describes the comparisons using 0 as the threshold. It is evident that the standard FS test is a special case of the FS-MT test, by only involving the second column of comparisons for calculating the win-loss score.